OpenStudy (anonymous):

The area of a rectangle is 96 square centimeters. If the width of the rectangle is 4 centimeters less than its length, what are the dimensions of the rectangle?

OpenStudy (anonymous):

width is given = 4 cm , length = area / width = 96 / 4=24 cm

OpenStudy (anonymous):

let length = x cm then width = (x-4) cm area of rect = l x b so x (x-4) = 96 x^2 - 4x - 96 = 0 x^2 -12x +8x - 96 = 0 x(x-12) +8(x-12) = 0 (x+8)(x-12) = 0 this is possible only if either x+8 = 0 or x-12 = 0 if x+8 = 0 then x= -8 but this is not possible as length cannot be negative if x-12 = 0 then x = 12 So length of rectangle is 12 cm and breadth is 12-4 = 8cm

OpenStudy (anonymous):

if u r satisfied with the answer, pls click the GOOD ANSWER button to show appreciation and to mark the question as solved and closed